OpenStudy (anonymous):

2x-y=3, 7x-3y=12 substution method

OpenStudy (akashdeepdeb):

@uri You aren't substituting there! :)

OpenStudy (uri):

lol

OpenStudy (anonymous):

would y be 3 as well?

OpenStudy (uri):

@terenzreignz Am i wrong..

terenzreignz (terenzreignz):

No... but the instructions *clearly* say 'use substitution'...

OpenStudy (akashdeepdeb):

You eliminated there! :)

OpenStudy (uri):

Ow.

terenzreignz (terenzreignz):

Try it again @uri ^_^

OpenStudy (uri):

Le doing accounting homework. :p

terenzreignz (terenzreignz):

You and your excuses... Anyway, @drea83 Your first step is to pick an equation (first one recommended) and isolate one of the variables (y recommended).

OpenStudy (anonymous):

uri , can you write the question with system ?

OpenStudy (anonymous):

Do you know system ?

OpenStudy (anonymous):

When we have x & y use that

OpenStudy (anonymous):

Look :!

OpenStudy (akashdeepdeb):

Do this... FOR SUBSTITUTION STEP 1: Find the value of any variable in terms of the other variable in the first equation. STEP 2: 2x-y=3 y = 2x - 3 STEP 3: NOW substitute this variable's value in the second equation. STEP 4: 7x-3y=12 7x - 3(2x-3) = 12 7x - 6x + 9 = 12 Get x's value. ^_^ [Learnt that from @terenzreignz ] STEP 5: Put x's value in the first equation to get the y value. :D Understood? :)

OpenStudy (nirmalnema):

2x-3=y and 7x-3y=12 so u get y=2x-3 from the 1st eq.. replace its value in the 2nd eq.. so u get 12=7x-3(2x-3) 12=7x-6x+9 x = 3 put in the 1st eq u will get y = 3

OpenStudy (akashdeepdeb):

@drea83 Understood this? :D

OpenStudy (anonymous):

2x-y=3 7x-3y=12 ___________________________________________________________________________________ 7 . (2x -y = 3) 2 . (7x-3y=12) _____________________________________________________________________________________- And now we want find y s first so shold x1= -(x2) ______________________________________________________________________________ so we have : 14x - 7y =21 14x - 6y =24 _________________________________________________________________________________ Now we should kill x s so reduse them : but we dont have : x1= -(x2) so do it by (-1) ___________________________________________________________________________________-- -1 .( 14x - 7y =21) ___________________________________________________________________________________ so we have : -14x + 7y =-21 14x-6y=+12 ______________________________ then reduse them from them ! -14x+7y =-21 14x -6y =12 __________________ y =-9 ________________________________________________________ Y = -9 Now can you find X ? Do you under stand ?